# DisCO: Reinforcing Large Reasoning Models with Discriminative Constrained Optimization

### v2

@@ Abstract
The recent success and openness of DeepSeek-R1 have brought widespread attention to Group Relative Policy Optimization (GRPO) as a reinforcement learning method for large reasoning models (LRMs). In this work, we analyze the GRPO objective under a binary reward setting and reveal an inherent limitation of question-level difficulty bias arising from its group relative advantage function. We also identify a connection between GRPO and traditional discriminative methods in supervised learning. Motivated by these insights, we introduce a new Discriminative Constrained Optimization (DisCO) framework for reinforcing LRMs, grounded in the principle of discriminative learning: increasing the scores of positive answers while decreasing those of negative ones. The main differences between DisCO and GRPO and its recent variants are: (1) it replaces the group relative objective with a discriminative objective defined by a scoring function; (2) it abandons clipping-based surrogates in favor of non-clipping RL surrogate objectives used as scoring functions; (3) it employs a simple yet effective constrained optimization approach to enforce the KL divergence constraint, ensuring stable training. As a result, DisCO offers notable advantages over GRPO and its variants: (i) it completely eliminates difficulty bias by adopting discriminative objectives; (ii) it addresses the entropy instability in GRPO and its variants through the use of non-clipping scoring functions and a constrained optimization approach; (iii) it allows the incorporation of advanced discriminative learning techniques to address data imbalance, where a significant number of questions have more negative than positive generated answers during training. Our experiments on enhancing the mathematical reasoning capabilities of SFT-finetuned models show that DisCO significantly outperforms GRPO and its improved variants such as DAPO, achieving average gains of 7% over GRPO and 6% over DAPO across six benchmark tasks for an 1.5B model.
Date: 05/19/2025
Github: https://github.com/Optimization-AI/DisCO
@@ 1 Introduction
The recent success and openness of DeepSeek-R1 have sparked a surge of interest in large reasoning models (LRMs), particularly in the context of fine-tuning via reinforcement learning (RL) [23]. The core approach involves iteratively generating synthetic data using the reasoning model and applying a rule-based reward mechanism to label the outputs. These rewards are then used to update the policy, i.e., the reasoning model itself. Notably, this framework, featuring a novel policy optimization method called Group Relative Policy Optimization (GRPO), has enabled DeepSeek-R1 to achieve performance comparable to advanced proprietary LRMs such as OpenAI-o1 on many reasoning benchmarks. As a result, GRPO has rapidly become a focal point for advancing LRM capabilities, particularly in domains like mathematics and scientific reasoning.
Several efforts have sought to replicate the performance of DeepSeek-R1 or to further enhance reasoning models using GRPO [67, 31, 42, 26, 69, 74, 5], while few others have tried to identify its inherent limitations with potential remedies [40, 79, 39]. Useful techniques have been introduced to improve GRPO [40, 41, 79, 26, 13, 82]. For instance, DAPO [79] employs two distinct clipping hyperparameters to mitigate entropy collapse, encouraging exploration. Dr. GRPO [40] removes the variance normalization in advantage function, aiming to mitigate the issue of difficulty bias. However, these approaches remain heuristic and ad-hoc, lacking a principled foundation and falling short of fully addressing GRPO’s inherent limitations. Our analysis identifies that Dr. GRPO continues to suffer from the difficulty bias issue, while our experiments show that DAPO may induce excessive entropy growth, producing highly random outputs. This motivates us to explore a central question:
@@
Building upon these insights, we propose a principled optimization framework for reinforcing large reasoning models based on discriminative learning. Specifically, we optimize a discriminative objective using a proper scoring function over input-output pairs, which increases the score of positive outputs and decreases that of negative ones. The flexibility of our framework allows us to leverage simple non-clipping RL surrogate objectives as scoring functions without suffering from entropy instability, and to incorporate advanced discriminative techniques to address data imbalance in generated rollouts. To ensure training stability, we adopt a simple yet effective constrained optimization method to enforce a trust region constraint bounding the KL divergence between the updated model and the old model. Our experiments for mathematical reasoning show that DisCO significantly outperforms all baselines for fine-tuning DeepSeek-R1-Distill-Qwen-1.5B/7B models with a maximum 8k response length for both training and inference, and also achieves a better performance than GRPO that uses a maximum 24k length for training and 32k length for inference.
@@
  • We demonstrate significant improvements of our DisCO method over GRPO and six other baselines, including DAPO, through experiments for fine-tuning LRMs on mathematical reasoning tasks, with evaluations across six benchmarks.
@@ 2 Related Work
Large Reasoning Models (LRMs). Recent advances of LRMs, such as OpenAI o1 [51], DeepSeek-R1 [23] and Kimi K1.5 [63], have demonstrated strong reasoning capability in solving complex tasks. Departing from earlier approaches in LLMs, such as Chain-of-thought (CoT) prompting [66, 48, 81], Tree-of-Thought [78], Monte Carlo Tree Search [18, 64, 71], a major breakthrough was achieved by scaling RL training using verifiable rewards to incentivize LLMs to learn through self-exploration [23, 63]. Inspired by DeepSeek-R1’s core algorithm GRPO [59], the research community has actively pursued improved techniques for large-scale RL training, focusing primarily on three directions: algorithm design [79, 40, 13, 39, 63, 61], reward curation [82, 67, 79], and sampling strategies [79, 26, 83, 29]. Our work falls under the category of algorithm design.
@@
Among these, Dr. GRPO [40] identifies response-level length bias and question-level difficulty bias in GRPO algorithm, advocating the removal of length and advantage normalization to improve token efficiency. DAPO [79] highlights several limitations of GRPO, such as entropy collapse, training instability, and biased loss, and addresses them through techniques like decoupled clipping, dynamic sampling, and a token-level policy loss. GPG [13] introduces a simplified REINFORCE-based objective that eliminates the need for both the critic and reference models, thereby enhancing scalability for RL training. TRPA [61] simply uses the Direct Preference Optimization (DPO) objective and a KL divergence regularization for fine-tuning LRMs. It can be recovered from our basic approach, which uses a logistic function as the surrogate loss, the log of likelihood ratio with respect to a frozen reference model as the scoring function, and the KL divergence as a regularization rather than a constraint. However, it does not address the imbalanced rollouts. The uniqueness and significance of our contributions lie in the analysis of GRPO objective and its variants that reveal key limitations, and the integration of advanced discriminative learning approaches for handling imbalanced rollouts and efficient constrained optimization technique for ensuring training stability.
@@
Reinforcement Learning (RL). RL is a learning paradigm centered on control and decision-making, in which an agent optimizes a target objective through trial-and-error interactions with its environment [8]. RL approaches are typically categorized into model-based [60, 53, 49, 17] and model-free methods [68, 62, 46, 56, 57, 38, 20]. Among model-free methods, the evolution from Vanilla Policy Gradient [68, 62] to TRPO [56] and PPO [57] has influenced the development of GRPO. In the context of fine-tuning LLMs, another line of work is RL from human feedback (RLHF). An early example of connecting RL with LLMs dates back to OpenAI’s work on integrating human preferences to improve text generation tasks, such as summarization using the PPO algorithm [85]. This approach was later extended to fine-tune LLMs for instruction following and/or alignment on helpfulness and harmlessness [52, 3, 22, 75]. Due to the high data requirements and training costs of standard RLHF, off-policy methods like DPO [54] and its variants [2, 16, 72, 44, 24], have been proposed to reduce reliance on explicit reward models. Another line of on-policy algorithms for RLHF, such as RLOO [1], ReMax [36], and REINFORCE++ [28], has been introduced to reduce the computational burden by removing the critic network in PPO. While some works [43, 32, 28, 10, 70] attempt to adapt RLHF techniques for reasoning tasks, they have not yielded significant improvements.
@@
Discriminative Learning. Parallel to RL, discriminative learning is another classical learning paradigm, that has been studied extensively for many traditional tasks, including multi-class classification [14, 15, 6], AUC maximization [76, 80], and learning to rank [9, 19, 7]. These methods are grounded in the common principle of increasing prediction scores for positive (relevant) labels (data) while decreasing scores for negative (irrelevant) ones. Nevertheless, discriminative learning remains under-explored in the cotext of LLM training. Recently, Guo et al. [25] proposed discriminative probabilistic approaches for supervised fine-tuning of LLMs. However, unlike our approach, they did not employ an RL framework with verifiable rewards to fine-tune LRMs.
@@ 3 Preliminaries
We consider fine-tuning a generative reasoning model $\pi_{\theta}$ parameterized by $\theta$. The old model in one step of learning is denoted by $\pi_{\text{old}}$. It is used to generate answers for a set of input questions. Given a question $q$ (with prompt included), the generated output $o$ follows the distribution $\pi_{\text{old}}(\cdot|q)$, which includes reasoning traces and the final answer. Specifically, output $o$ is generated token by token, i.e., $o_{t}\sim\pi_{\text{old}}(\cdot|q,o_{<t})$, for $t=1,\cdots,|o|$. We consider a rule-based reward mechanism that returns a binary value for a given question $q$ and its corresponding answer in the output $o$, which uses either exact match against extracted answer or a formal verification tool [23, 33, 55]. Let $r(o|q)\in\{1,0\}$ denote the reward assigned to an output $o$ with respect to the input $q$. Let $p(q)=\mathbb{E}_{o\sim\pi_{\text{old}}(\cdot|q)}[r(o|q)]\in[0,1]$, which quantifies the difficulty of the question $q$ under the model $\pi_{\text{old}}$. We denote by $\pi_{\text{old}}^{+}(\cdot|q)$ the conditional distribution of outputs when the reward is one (i.e., positive answers) and by $\pi_{\text{old}}^{-}(\cdot|q)$ the conditional distribution of outputs when the reward is zero (i.e., negative answers). By the law of total expectation, for any function $g(o,q)$ we have
@@
Recently, several variants of GRPO have been introduced [79, 13, 39, 82, 40]. Many of them retain the advantage function $A(o|q)$ while modifying other components such as hyper-parameter $\epsilon$, the normalization factor and the likelihood ratio. Several works employ an unnormalized advantage function $\hat{A}(o|q)=r(o|q)-\mathbb{E}_{o^{\prime}\sim\pi_{\text{old}}(\cdot|q)}r(o^{\prime}|q)$ [40, 13].
@@ 4 Analysis of GRPO and its variants
In the following analysis we assume $p(q)\in(0,1)$; otherwise we can remove them from consideration as done in practice [23, 42, 59].
@@
Assume that $f(x,y)$ is non-decreasing function of $x$ such that $f(x,y)=\mathbb{I}(y>0)yf^{+}(x,1)-\mathbb{I}(y\leq 0)yf^{-}(x,1)$, where both $f^{+},f^{-}$ are non-decreasing functions of $x$, then we have
@@
$s_{\theta}^{+}(o,q)$ $=\frac{1}{|o|}\sum_{t=1}^{|o|}\min(\frac{\pi_{\theta}(o_{t}|q,o_{<t})}{\pi_{\text{old}}(o_{t}|q,o_{<t})},1+\epsilon),\quad s_{\theta}^{-}(o,q)=\frac{1}{|o|}\sum_{t=1}^{|o|}\max(\frac{\pi_{\theta}(o_{t}|q,o_{<t})}{\pi_{\text{old}}(o_{t}|q,o_{<t})},1-\epsilon).$ (5)
@@
The proof of the above proposition is included in Appendix B.1 and is inspired by [47] with differences that lead to two new insights from Proposition 1 regarding the two components of $\mathcal{J}_{0}$. First, let us consider the component $\mathbb{E}_{o\sim\pi_{\text{old}}^{+}(\cdot|q),o^{\prime}\sim\pi_{\text{old}}^{-}(\cdot|q),}[s_{\theta}^{+}(o,q)-s_{\theta}^{-}(o^{\prime},q)]$. Since both $f^{+}$ and $f^{-}$ are non-decreasing functions of the first argument, then both $s_{\theta}^{+}(o,q)$ and $s_{\theta}^{-}(o,q)$ are non-decreasing functions of $\pi_{\theta}(o_{t}|q,o_{<t})$. Hence, maximizing $\mathcal{J}_{0}$ would increase the likelihood of tokens in the positive answers and decrease the likelihood of tokens in the negative answers. This makes sense as we would like the new model to have a high likelihood of generating a positive (correct) answer and a low likelihood of generating a negative (incorrect) answer. This mechanism is closely related to traditional discriminative methods of supervised learning in the context of AUC maximization [77], which aims to maximize the scores of positive samples $o\sim\pi_{\text{old}}^{+}(\cdot|q)$ while minimizing scores of negative samples $o^{\prime}\sim\pi_{\text{old}}^{-}(\cdot|q)$, where the $q$ acts like the classification task in the AUC maximization. Hence, in the context of discriminative learning, we refer to $s^{+}(o,q)$ and $s^{-}(o,q)$ as scoring functions. Therefore, $\mathbb{E}_{o\sim\pi_{\text{old}}^{+}(\cdot|q),o^{\prime}\sim\pi_{\text{old}}^{-}(\cdot|q),}[s^{+}(o,q)-s^{-}(o^{\prime},q)]$ is a discriminative objective.
@@
Second, let us consider the component $\omega(q)=\sqrt{p(q)(1-p(q))}$, which acts like a weight scaling the discriminative objective for each individual input question. It is this component that leads to difficulty bias. As shown in Figure LABEL:fig:intro(a), questions with very high $p(q)$ values (close to 1) or very low $p(q)$ values (close to 0) receive small weights for their discriminative objectives, causing the optimization to focus primarily on questions of intermediate difficulty while paying little attention to hard questions ($p(q)\approx 0$) and easy questions ($p(q)\approx 1$). This mechanism may significantly hinder the learning efficiency. Intuitively, if the generated answers have only one correct solution out of 10 trials, i.e. $p(q)=0.1$, we should grasp this chance to enhance the model instead of overlooking it. On the other hand, even when we encounter an easy question with a probability of $p(q)=0.9$, we should keep improving the model rather than being satisfied because it still makes mistakes with respect to this question. Our hypothesis is that removing this weight could accelerate the training. To validate this hypothesis, we conducted a series of empirical experiments for fine-tuning an 1.5B model as described in Section 6. We start by examining whether a substantial number of questions have correctness probabilities ($p(q)$) near 0 or 1. As shown in Figure LABEL:fig:intro(b), during GRPO training, the correctness probabilities across individual questions appear broadly distributed, with many near 0 or 1. Then, we compare the original GRPO with a variant below without the weight $\sqrt{p(q)(1-p(q))}$:
@@
The results are shown in Figure LABEL:fig:intro(c) and  LABEL:fig:intro(d). We can observe that the variant without the weighting mechanism quickly achieves a higher ratio of 100% correctness and a lower ratio of 0% correctness, confirming the detrimental impact of the inappropriate weighting.
We note that the difficulty bias has been pointed out in a recent work Dr. GRPO [40]. To mitigate this issue, Dr. GRPO uses the un-normalized advantage function $\hat{A}(o|q)$. However, with a similar analysis as above (cf. Appendix B.3), we can derive that Dr. GRPO still has a question-level weight $\omega(q)=p(q)(1-p(q))$ before the discriminative objective. As shown in Figure LABEL:fig:intro(a), this weight mitigates but does not eliminate the imbalanced weight across questions.

### v4

@@ Abstract
The recent success and openness of DeepSeek-R1 have brought widespread attention to Group Relative Policy Optimization (GRPO) as a reinforcement learning method for large reasoning models (LRMs). In this work, we analyze the GRPO objective under a binary reward setting and reveal an inherent limitation of question-level difficulty bias arising from its group relative advantage function. We also identify a connection between GRPO and traditional discriminative methods in supervised learning. Motivated by these insights, we introduce a new Discriminative Constrained Optimization (DisCO) framework for reinforcing LRMs, grounded in the principle of discriminative learning: increasing the scores of positive answers while decreasing those of negative ones. The main differences between DisCO and GRPO and its recent variants are: (1) it replaces the group relative objective with a discriminative objective defined by a scoring function; (2) it abandons clipping-based surrogates in favor of non-clipping RL surrogate objectives used as scoring functions; (3) it employs a simple yet effective constrained optimization approach to enforce the KL divergence constraint. As a result, DisCO offers notable advantages over GRPO and its variants: (i) it completely eliminates difficulty bias by adopting discriminative objectives; (ii) it addresses the entropy instability in GRPO and its variants through the use of non-clipping scoring functions and a constrained optimization approach, yielding long and stable training dynamics; (iii) it allows the incorporation of advanced discriminative learning techniques to address data imbalance, where a significant number of questions have more negative than positive generated answers during training. Our experiments on enhancing the mathematical reasoning capabilities of SFT-finetuned models show that DisCO significantly outperforms GRPO and its improved variants such as DAPO, achieving average gains of 7% over GRPO and 6% over DAPO across six benchmark tasks for an 1.5B model. [^1] The code is available at: https://github.com/Optimization-AI/DisCO
@@ 1 Introduction
The recent success and openness of DeepSeek-R1 have sparked a surge of interest in large reasoning models (LRMs), particularly in the context of fine-tuning via reinforcement learning (RL) guo2025deepseek. The core approach involves iteratively generating synthetic data using the reasoning model and applying a rule-based reward mechanism to label the outputs. These rewards are then used to update the policy, i.e., the reasoning model itself. Notably, this framework, featuring a novel policy optimization method called Group Relative Policy Optimization (GRPO), has enabled DeepSeek-R1 to achieve performance comparable to advanced proprietary LRMs at that time such as OpenAI-o1 on many reasoning benchmarks. As a result, GRPO has rapidly become a focal point for advancing LRM capabilities, particularly in domains like mathematics and scientific reasoning.
Several efforts have sought to replicate the performance of DeepSeek-R1 or to further enhance reasoning models using GRPO wen2025light; openr1; deepscaler2025; skywork-or1-2025; xiaomi2025mimo; xue2025dancegrpounleashinggrpovisual; bercovich2025llamanemotronefficientreasoningmodels, while few others have tried to identify its inherent limitations with potential remedies liu2025understanding; yu2025dapo; lin2025cppo. Useful tricks have been introduced to improve GRPO liu2025understanding; deepcoder2025; yu2025dapo; skywork-or1-2025; chu2025gpg; zhang2025grpo. For instance, DAPO yu2025dapo employs two distinct clipping hyperparameters to mitigate entropy collapse, encouraging exploration. Dr. GRPO liu2025understanding removes the variance normalization in advantage function, aiming to mitigate the issue of difficulty bias. However, these approaches remain heuristic and ad-hoc, lacking a principled foundation and falling short of fully addressing GRPO’s inherent limitations. Our analysis identifies that Dr. GRPO continues to suffer from the difficulty bias issue, while our experiments show that DAPO may induce excessive entropy growth, producing highly random outputs. This motivates us to explore a central question:
@@
Building upon these insights, we propose a principled optimization framework for reinforcing large reasoning models based on discriminative learning. Specifically, we optimize a discriminative objective using a proper scoring function over input-output pairs, which increases the score of positive outputs and decreases that of negative ones. The flexibility of our framework allows us to leverage simple non-clipping RL surrogate objectives as scoring functions without suffering from entropy instability, and to incorporate advanced discriminative techniques to address data imbalance in generated rollouts. To ensure training stability, we adopt a simple yet effective constrained optimization method to enforce a trust region constraint bounding the KL divergence between the updated model and the old model. Our experiments for mathematical reasoning show that DisCO significantly outperforms all baselines for fine-tuning DeepSeek-R1-Distill-Qwen and -Llama models with a maximum 8k response length for both training and inference, and also achieves a better performance than GRPO that uses a maximum 24k length for training and 32k length for inference.
@@
  • We demonstrate significant improvements of our DisCO method over GRPO and four other baselines, including DAPO, through experiments for fine-tuning LRMs on mathematical reasoning tasks, with evaluations across six benchmarks.
@@ 2 Related Work
Large Reasoning Models (LRMs). Recent advances of LRMs, such as OpenAI o1 openo1, DeepSeek-R1 guo2025deepseek and Kimi K1.5 team2025kimi, have demonstrated strong reasoning capability in solving complex tasks. Departing from earlier approaches in LLMs, such as Chain-of-thought (CoT) prompting wei2022chain; muennighoff2025s1; zelikman2022star, Tree-of-Thought yao2023tree, Monte Carlo Tree Search feng2023alphazero; trinh2024solving; xin2024deepseek, a major breakthrough was achieved by scaling RL training using verifiable rewards to incentivize LLMs to learn through self-exploration guo2025deepseek; team2025kimi. Inspired by DeepSeek-R1’s core algorithm GRPO shao2024deepseekmath, the research community has actively pursued improved techniques for large-scale RL training, focusing primarily on three directions: algorithm design yu2025dapo; liu2025understanding; chu2025gpg; lin2025cppo; team2025kimi; su2025trust, reward curation zhang2025grpo; wen2025light; yu2025dapo, and sampling strategies yu2025dapo; skywork-or1-2025; zhang2025srpo; hu2025open. Our work falls under the category of algorithm design.
@@
Among these, Dr. GRPO liu2025understanding identifies response-level length bias and question-level difficulty bias in GRPO algorithm, advocating the removal of length and advantage normalization to improve token efficiency. DAPO yu2025dapo highlights several limitations of GRPO, such as entropy collapse, training instability, and biased loss, and addresses them through techniques like decoupled clipping, dynamic sampling, and a token-level policy loss. GPG chu2025gpg introduces a simplified REINFORCE-based objective that eliminates the need for both the critic and reference models, thereby enhancing scalability for RL training. TRPA su2025trust simply uses the Direct Preference Optimization (DPO) objective and a KL divergence regularization for fine-tuning LRMs. It can be recovered from our basic approach, which uses a logistic function as the surrogate loss, the log of likelihood ratio with respect to a frozen reference model as the scoring function, and the KL divergence as a regularization rather than a constraint. However, it does not address the imbalanced rollouts. The uniqueness and significance of our contributions lie in the analysis of GRPO objective and its variants that reveal key limitations, and the integration of advanced discriminative learning approaches for handling imbalanced rollouts and efficient constrained optimization technique for ensuring training stability.
@@
Reinforcement Learning (RL). RL is a learning paradigm centered on control and decision-making, in which an agent optimizes a target objective through trial-and-error interactions with its environment cao2024survey. RL approaches are typically categorized into model-based silver2017mastering; racaniere2017imagination; nagabandi2018neural; feinberg2018model and model-free methods williams1992simple; sutton1999policy; mnih2016asynchronous; schulman2015trust; schulman2017proximal; lillicrap2015continuous; fujimoto2018addressing. Among model-free methods, the evolution from Vanilla Policy Gradient williams1992simple; sutton1999policy to TRPO schulman2015trust and PPO schulman2017proximal has influenced the development of GRPO. In the context of fine-tuning LLMs, another line of work is RL from human feedback (RLHF). An early example of connecting RL with LLMs dates back to OpenAI’s work on integrating human preferences to improve text generation tasks, such as summarization using the PPO algorithm ziegler2019fine. This approach was later extended to fine-tune LLMs for instruction following and/or alignment on helpfulness and harmlessness ouyang2022training; bai2022training; grattafiori2024llama; yang2024qwen2. Due to the high data requirements and training costs of standard RLHF, off-policy methods like DPO rafailov2023direct and its variants azar2024general; ethayarajh2024kto; xu2024contrastive; meng2024simpo; guo2025discriminative, have been proposed to reduce reliance on explicit reward models. Another line of on-policy algorithms for RLHF, such as RLOO ahmadian2024back, ReMax li2023remax, and REINFORCE++ hu2025reinforce, has been introduced to reduce the computational burden by removing the critic network in PPO. While some works luong2024reft; kazemnejad2024vineppo; hu2025reinforce; chen2024step; xie2024monte attempt to adapt RLHF techniques for reasoning tasks, they have not yielded significant improvements.
@@
Discriminative Learning. Parallel to RL, discriminative learning is another classical learning paradigm, that has been studied extensively for many traditional tasks, including multi-class classification cortes1995support; 10.5555/944790.944813; bishop2006pattern, AUC maximization yang2022aucmaximizationerabig; yuan2021large, and learning to rank listnet; freund2003efficient; burges2005learning. These methods are grounded in the common principle of increasing prediction scores for positive (relevant) labels (data) while decreasing scores for negative (irrelevant) ones. Nevertheless, discriminative learning remains under-explored in the cotext of LLM training. Recently, Guo et al. guo2025discriminativefinetuninggenerativelarge proposed discriminative probabilistic approaches for supervised fine-tuning of LLMs. However, unlike our approach, they did not employ an RL framework with verifiable rewards to fine-tune LRMs.
@@ 3 Preliminaries
We consider fine-tuning a generative reasoning model $\pi_{\theta}$ parameterized by $\theta$. The old model in one step of learning is denoted by $\pi_{\text{old}}$. It is used to generate answers for a set of input questions. Given a question $q$ (with prompt included), the generated output $o$ follows the distribution $\pi_{\text{old}}(\cdot|q)$, which includes reasoning traces and the final answer. Specifically, output $o$ is generated token by token, i.e., $o_{t}\sim\pi_{\text{old}}(\cdot|q,o_{<t})$, for $t=1,\cdots,|o|$. We consider a rule-based reward mechanism that returns a binary value for a given question $q$ and its corresponding answer in the output $o$, which uses either exact match against extracted answer or a formal verification tool guo2025deepseek; lambert2025tulu3pushingfrontiers; ren2025deepseekproverv2advancingformalmathematical. Let $r(o|q)\in\{1,0\}$ denote the reward assigned to an output $o$ with respect to the input $q$. Let $p(q)=\mathbb{E}_{o\sim\pi_{\text{old}}(\cdot|q)}[r(o|q)]\in[0,1]$, which quantifies the difficulty of the question $q$ under the model $\pi_{\text{old}}$. We denote by $\pi_{\text{old}}^{+}(\cdot|q)$ the conditional distribution of outputs when the reward is one (i.e., positive answers) and by $\pi_{\text{old}}^{-}(\cdot|q)$ the conditional distribution of outputs when the reward is zero (i.e., negative answers). By the law of total expectation, for any function $g(o,q)$ we have
@@
Recently, several variants of GRPO have been introduced yu2025dapo; chu2025gpg; lin2025cppo; zhang2025grpo; liu2025understanding. Many of them retain the advantage function $A(o|q)$ while modifying other components such as hyper-parameter $\epsilon$, the normalization factor and the likelihood ratio. Several works employ an unnormalized advantage function $\hat{A}(o|q)=r(o|q)-\mathbb{E}_{o^{\prime}\sim\pi_{\text{old}}(\cdot|q)}r(o^{\prime}|q)$ liu2025understanding; chu2025gpg.
@@ 4 Analysis of GRPO and its variants
In the following analysis we assume $p(q)\in(0,1)$; otherwise we can remove them from consideration as done in practice guo2025deepseek; deepscaler2025; shao2024deepseekmath.
@@
Assume that $f(x,y)$ is non-decreasing function of $x$ such that $f(x,y)=\mathbb{I}(y>0)yf^{+}(x,1)-\mathbb{I}(y\leq 0)|y|f^{-}(x,1)$, where both $f^{+},f^{-}$ are non-decreasing functions of $x$, then we have
@@
$f^{+}(x,1)=\min(x,1+\epsilon),\quad f^{-}(x,1)=\max(x,1-\epsilon).$ (5)
@@
The proof of the above proposition is included in Appendix B.1 and is inspired by mroueh2025reinforcementlearningverifiablerewards with differences that lead to two new insights from Proposition 1 regarding the two components of $\mathcal{J}_{0}$. First, let us consider the component $\mathbb{E}_{o\sim\pi_{\text{old}}^{+}(\cdot|q),o^{\prime}\sim\pi_{\text{old}}^{-}(\cdot|q)}[s_{\theta}^{+}(o,q)-s_{\theta}^{-}(o^{\prime},q)]$. Since both $f^{+}$ and $f^{-}$ are non-decreasing functions of the first argument, then both $s_{\theta}^{+}(o,q)$ and $s_{\theta}^{-}(o,q)$ are non-decreasing functions of $\pi_{\theta}(o_{t}|q,o_{<t})$. Hence, maximizing $\mathcal{J}_{0}$ would increase the likelihood of tokens in the positive answers and decrease the likelihood of tokens in the negative answers. This makes sense as we would like the new model to have a high likelihood of generating a positive (correct) answer and a low likelihood of generating a negative (incorrect) answer. This mechanism is closely related to traditional discriminative methods of supervised learning in the context of AUC maximization yang2022auc, which aims to maximize the scores of positive samples $o\sim\pi_{\text{old}}^{+}(\cdot|q)$ while minimizing scores of negative samples $o^{\prime}\sim\pi_{\text{old}}^{-}(\cdot|q)$, where the $q$ acts like the classification task in the AUC maximization. Hence, in the context of discriminative learning, we refer to $s^{+}(o,q)$ and $s^{-}(o,q)$ as scoring functions. Therefore, $\mathbb{E}_{o\sim\pi_{\text{old}}^{+}(\cdot|q),o^{\prime}\sim\pi_{\text{old}}^{-}(\cdot|q),}[s^{+}(o,q)-s^{-}(o^{\prime},q)]$ is a discriminative objective.
(a) (a) Weight on questions based on correctness probability $p$; (b) Histogram of per-question accuracy evaluated in the GRPO learning; (c) Comparison of the ratio of questions with 100% correctness probability; (d) Comparison of the ratio of questions with 0% correctness probability.
@@
Second, let us consider the component $\omega(q)=\sqrt{p(q)(1-p(q))}$, which acts like a weight scaling the discriminative objective for each individual input question. It is this component that leads to difficulty bias. As shown in Figure 4(a)(a), questions with very high $p(q)$ values (close to 1) or very low $p(q)$ values (close to 0) receive small weights for their discriminative objectives, causing the optimization to focus primarily on questions of intermediate difficulty while paying little attention to hard questions ($p(q)\approx 0$) and easy questions ($p(q)\approx 1$). This mechanism may significantly hinder the learning efficiency. Intuitively, if the generated answers have only one correct solution out of 10 trials, i.e. $p(q)=0.1$, we should grasp this chance to enhance the model instead of overlooking it. On the other hand, even when we encounter an easy question with a probability of $p(q)=0.9$, we should keep improving the model rather than being satisfied because it still makes mistakes with respect to this question. Our hypothesis is that removing this weight could accelerate the training. To validate this hypothesis, we conducted a series of empirical experiments for fine-tuning a 1.5B model as described in Section 6. We start by examining whether a substantial number of questions have correctness probabilities ($p(q)$) near 0 or 1. As shown in Figure 4(a)(b), during GRPO training, the correctness probabilities across individual questions appear broadly distributed, with many near 0 or 1. Then, we compare the original GRPO with a variant that removes weight $\sqrt{p(q)(1-p(q))}$:
@@
The results are shown in Figure 4(a)(c) and  4(a)(d). We can observe that the variant without the weighting mechanism quickly achieves a higher ratio of 100% correctness and a lower ratio of 0% correctness, confirming the detrimental impact of the inappropriate weighting.
We note that the difficulty bias has been pointed out in a recent work Dr. GRPO liu2025understanding. To mitigate this issue, Dr. GRPO uses the un-normalized advantage function $\hat{A}(o|q)$. However, with a similar analysis as above (cf. Appendix B.3), we can derive that Dr. GRPO still has a question-level weight $\omega(q)=p(q)(1-p(q))$ before the discriminative objective. As shown in Figure 4(a)(a), this weight mitigates but does not eliminate the imbalanced weight across questions.